OpenStudy (anonymous):

integrate dy/dx = (y^2-1)/(x^2 - 1) my solutions manual says that simplifies to .... (1/2)(1/(y-1) - 1/(y+1)) dy= (1/2)(1/(x-1) - 1/(x+1)) dx how do they reach this answer?

OpenStudy (solomonzelman):

\(\large\color{slate}{ dy/dx = (y^2-1)/(x^2 - 1)}\) dividing both sides by y^2-1, and multiplying times dx on both sides: \(\large\color{slate}{ dy/(y^2-1) =dx/(x^2-1) }\) then both sides go by partical fraction.

OpenStudy (anonymous):

Partial fractions

OpenStudy (anonymous):

That's the short answer. Here's a more detailed explanation: The goal is to be able to write a rational function \(R(x)=\dfrac{P(x)}{Q(x)}\) as a sum of rational functions made up of the irreducible factors of the denominator, \(Q(x)\). In this case, you have \[R(x)=\frac{1}{x^2-1}=\frac{1}{(x-1)(x+1)}\] We have two linear factors, so we suppose we can split up \(R(x)\) into the following: \[\frac{1}{(x-1)(x+1)}=\frac{a_1}{x-1}+\frac{a_2}{x+1}\] for some real \(a_1,a_2\). If we distribute the denominator on the LHS to the RHS and cancel the relevant factors, we have \[1=a_1(x+1)+a_2(x-1)~~\iff~~0x+1=(a_1+a_2)x+(a_1-a_2)\] Matching up coefficients for corresponding powers of \(x\), we have \[\begin{cases}a_1+a_2=0\\a_1-a_2=1\end{cases}\]

OpenStudy (anonymous):

Thank you so much!